Tyrion:

Sonia works at a bakery. The function f(x) represents the amount of money Sonia earns per loaf, where x is the number of loaves she makes. The function g(x) represents the number of bread loaves Sonia bakes per hour, where x is the number of hours she works. Show all work to find f(g(x)), and explain what f(g(x)) represents. f(x) = 9x2 + 1 g of x equals the square root of the quantity 2 x cubed

Angle:

\(f(x) = 9x^2 + 1\) \(g(x) = \sqrt{2x^3}\) \(f(g(x)) = 9(~g(x)~)^2+1\) \(= 9(\sqrt{2x^3})^2+1\)

Tyrion:

thats how it needs to be set up??

Angle:

yes

Tyrion:

thank u